Institut for Matematik og Datalogi Odense Universitet

April 7, 1995 JFB

DM15 – F95 – Week 25

Lecture, April 3

We finished chapter 1 in Mitchell.

Lecture, April 10

We will begin on chapter 2 in Mitchell’s monograph.

Lecture, April 24

We will continue on the second chapter of Mitchell’s monograph.

Problems to be discussed in week 17

• Do exercises 1.3.1 and 1.3.2 on page 22 of Mitchell.

• Do exercises 1.6.1 on page 29 of Mitchell. Part (b) is obviously incorrect. Correct it before solving the problem.

• Do exercise 1.6.2 on page 29 of Mitchell. Define the functions fst, snd, and trd for extracting the different components of A_{1} × (A_{2} × A_{3}).

• Do exercises 2.2.2 and 2.2.3 on page 54 of Mitchell.

Clarifications of statements in the textbook

• Note that the sizes given for the power sets at the bottom of page 28 are incorrect. One of the exercises above asks you to correct them.